Let the sample space be S={1, 2, 3, 4, 5, 6, 7, 8, 9, 10}. Suppose the outcomes are equally likely. Compute the probability of the event E=​"an even number less than 7​."




1
Expert's answer
2022-02-21T18:06:38-0500

The probability of each outcome is "\frac{1}{10}=0.1."

There are three even numbers less than 7.

P(X=an even number less than 7)"=3*0.1=0.3."